# The fundamental localization phases in quasiperiodic systems: A unified framework and exact results

### v2

@@ II.3 Renormalization group
The renormalization group (RG) approach (Suslov, 1982; Thouless and Niu, 1983; Ostlund and Pandit, 1984; Niu and Nori, 1986, 1990; Szabó and Schneider, 2018; Jagannathan, 2021; Gonçalves et al., 2022, 2023b) investigates the relevance of renormalized coefficients by iterating the rational approximations of the QP parameter. Through this process, the RG determines the asymptotic behavior of the target eigenstates. With the rational approximation, the approximated Hamiltonian exhibits a periodic structure, where the energy dispersion becomes a periodic function of quasi-momentum $\kappa_{x}$ and $\kappa_{y}$. To systematically investigate how the dispersion coefficients evolve as the system size increases, we compute the characteristic determinant $P^{(n)}=|{\cal H}^{(n)}-E|$ at a target energy $E$ for a given $L=F_{n}$, which is given by
@@
where $\epsilon_{R}^{(n)}$ represents higher harmonic terms. The asymptotic nature of eigenstates at $n\rightarrow\infty$ is determined by comparing the relative magnitudes of renormalized coefficients. For extended states, the dominant hopping is along $x$ direction, yielding $\big{|}\mu_{R}^{(n)}/t_{R}^{(n)}\big{|},\big{|}V_{R}^{(n)}/t_{R}^{(n)}\big{|}\rightarrow 0$. For localized states, the dominant hopping is along $y$ direction, leading to $\big{|}t_{R}^{(n)}/V_{R}^{(n)}\big{|},\big{|}\mu_{R}^{(n)}/V_{R}^{(n)}\big{|}\rightarrow 0$. For critical states, all the three coefficients reamin comparable, such that $\big{|}\mu_{R}^{(n)}/V_{R}^{(n)}\big{|},\big{|}\mu_{R}^{(n)}/t_{R}^{(n)}\big{|}\geq 1$. The corresponding transition points between extend, localized and critical can be obtained accordingly, and the mobility edges corresponds to the energy dependent transition points.
@@ II.4 Avila’s global theory
When the Hamiltonian in Eq. (1) can be reduced to an effectively spinless quasiperiodic (QP) chain of dressed particles, the system can be analytically characterized using Avila’s global theory Avila (2015). This theory enables the analytical calculation of the Lyapunov exponent (LE), which is the inverse of the localization length. The method provides a rigorous characterization of extended, localized, and critical states. Specifically, the non-negative LE $\gamma(E)$, quantifies the localization properties of eigenstates with energy $E$. If $\gamma(E)>0$, the state is localized, with the localization length $\xi(E)=\gamma^{-1}(E)$. If $\gamma(E)=0$, the state is delocalized, with a divergent localization length, corresponding to either an extended or a critical state. Critical states arise when the transfer matrix $T_{j}$ of the delocalized states becomes singular Simon and Spencer (1989); Jitomirskaya and Marx (2012). This singularity originates from the presence of incommensurately distributed zeros (IDZs) in the hopping coefficients Zhou et al. (2023), or from an unbounded QP potential, which effectively divides the system into infinite segments, acting as IDZs.
Furthermore, Avila’s global theory can also be employed to derive the analytical expression for the correlation length of extended states. This is achieved by performing the dual transformation of the Hamiltonian into dual space, followed by calculating the localization length of localized states in dual space, which is equivalent to the correlation length of the corresponding extended states in real space.
In the following, we outline the procedure for obtaining the LE using Avila’s global theory. For the generic 1D spinful QP Hamiltonian in Eq. (1), we consider the 4-by-4 transfer matrix $T_{j}$ given by $(\begin{array}[]{cccc}u_{j+1\uparrow},&u_{j+1\downarrow},&u_{j\uparrow},&u_{j\downarrow}\end{array})^{\intercal}=T_{j}(\begin{array}[]{cccc}u_{j\uparrow},&u_{j\downarrow},&u_{j-1\uparrow},&u_{j-1\downarrow}\end{array})^{\intercal}$, where $u_{j\sigma}$ is the wave function of eigenstate at site $j$ with spin $\sigma$. However, the 4-by-4 transfer matrix does not guarantee exact solvability. Exact solvability arises when the transfer matrix is reduced to a block-diagonal 2-by-2 form. When the system incorporates the local constraint introduced in Sec. III.3, the Hamiltonian simplifies to a 1D spinless QP model of dressed particles, featuring a single-frequency potential and nearest-neighbor hopping. Consequently, the effective transfer matrix becomes a 2-by-2 matrix. In this regime, the Lyapunov exponent $\gamma_{\epsilon}(E)$ is computed as
@@
Here, $E$ is the energy of the corresponding eigenstate, $||A||$ is the norm of the matrix A, i.e. the square root of the largest eigenvalue of $A^{\dagger}A$, ${\cal T}_{n,1}=\prod_{j=1}^{n}T_{j}$. The $\epsilon$ is the imaginary part of complexified quasiperiodic phase shift $\theta+i\epsilon$. The LE is generally challenging to compute analytically, even for the simple 2-by-2 transfer matrix $T_{j}$. Nevertheless, a key result from Avila’s global theory allows for the computation of the LE for all eigenstates: The LE $\gamma_{\epsilon}(E)$ is convex, continuous and piecewise linear with quantized right-derivative, given by
@@
The LE can be analytically obtained when the nonzero derivative is quantized to be exactly one integer for all range of $\epsilon$: $d\gamma_{\epsilon}/d\epsilon=2\pi\mathbb{Z}$. This condition implies the absence of turning points before which corresponds to there is no turning point before $\gamma_{\epsilon}$ intersects the $x-$ or $y-$axis, as illustrated in Fig. 1(b). In this case, the analytic expression for the LE is first obtained by calculating the complexified LE $\gamma_{\epsilon}$ at the limit $\epsilon\to\infty$, which is computationally straightforward. Using the properties that $\gamma_{\epsilon}(E)$ is convex, continuous, and piecewise linear with a quantized right-derivative, the analytic expression of the LE determined at $\epsilon\to\infty$ can be extended to all values of $\epsilon$. Finally, the exact LE is obtained by evaluating $\gamma_{\epsilon=0}$, which represents the original Lyapunov exponent that characterizes the properties of all eigenstates.
Figure 2: Three universal results for the spinful quasiperiodic chains. (a) Criteria for the system to exhibit pure phases without mobility edges. Here, $\lambda$ is the system parameter and $E$ denotes the energy. A and B refer to two distinct phases, which may correspond to extended, localized, or critical states. The system transitions to pure phases in the absence of mobility edges, which implies that the energy-dependent transition points $\lambda_{c}(E)$ become energy-independent, denoted as $\lambda_{c}$. (b) Mechanism for the emergence of critical states in the 1D spinful quasiperiodic (QP) system, generated by dual-invariant generalized incommensurate zeros (GIZs) in matrix elements, marked as the blue and red circles in real and dual spaces. (c) The local constraint that transforms the generic 1D spinful QP chain into a 1D spinless QP chain of dressed particles (represented by orange spheres), with energy-dependent or independent nearest-neighbor hopping coefficients $t_{j}^{\mathrm{eff}}(E)$ and on-site potentials $V_{j}^{\mathrm{eff}}(E)$. This transformation can be analytically tracked using Avila’s global theory.
@@ III Universal results for the quasiperiodic spinful chains
In this section, we show three universal results for the spinful quasiperiodic systems, which facilitates to establish the unified framework of all fundamental localization phases, as summarized in Fig. 2. First, we prove a generic condition under which mobility edges (MEs) disappear in the system, namely, the pure phases are obtained under this condition. Second, we uncover a new universal mechanism that critical states are protected by the generalized incommensurate zeros in matrix elements. Third, we outline the conditions, under which Avila’s global theory can be applied to obtain analytic solutions to the localization properties, correlation length, and the associated phase transition points and mobility edges.
@@ III.1 Criteria for pure phases without MEs
For the system to exhibit a pure phase without MEs, it must satisfy the following criteria: The on-site matrix $M_{j}$ are purely quasiperiodic, and the hopping coupling matrix $\Pi_{j}$ are either uniform or purely quasiperiodic, with the system satisfying the chiral symmetry
@@
Here $\Pi_{j}^{\prime}$ and $M_{j}^{\prime}$ are the transformed matrices obtained in the basis that diagonalizes the on-site coupling matrix, given explicitly as
@@
We demonstrate this criterion by proving that all leading coefficients in Eq.(2) are energy-independent, ensuring eigenstate transitions occur independently of energy, and consequently, the system exhibits pure phases without mobility edges (MEs) [Fig.2(a)]. This follows from a careful power counting analysis of the energy dependence in characteristic polynomial $P(E;\kappa_{x},\kappa_{y})$ in Eq. (2). We consider an odd system size $L$ without loss of generality. Due to chiral symmetry, eigenenergies appear in pairs $(-E,E)$, and therefore $P(E)$ is an even function of energy $E$. In the power counting analysis, the energy $E$ explicitly contributes a power of $E$, while $p_{j}^{\perp}$ and $p_{j}^{z}$ contribute dispersions of $e^{ik_{x}}$, and $m_{j}$ contributes dispersions of $e^{ik_{y}}$. Here $\kappa_{x}=Lk_{x}$ and $\kappa_{y}=Lk_{y}$. We utilize the periodic structure to simplify the power counting. Thus, the characteristic polynomial can be expressed as the determinant of a $2L\times 2L$ matrix $|A(E)|$ (details in Appendix A), with elements
@@
with $t_{j}^{\pm}=p_{j}^{z}\pm ip_{j}^{\perp}$. This structure ensures the energy terms and momentum dispersions decouple naturally, eliminating cross terms such as $Ee^{ik_{x}}$. Below, we separately analyze the dominant dispersions for extended, localized, and critical states.
@@
For the extended and localized states, characterized by dispersions of the form $\cos n_{x}Lk_{x}$ and $\cos n_{y}Lk_{y}$, where $n_{x},n_{y}\geq 1\in\mathbb{Z}$, we show that the dominated coefficients are energy-independent. Specifically, the leading dispersion $\cos Lk_{x}$ and $\cos Lk_{y}$ vanish due to chiral symmetry, as these terms arise from the product of $L$ factors of $e^{ik_{x}}$ and $L$ factors of energy $E$, yielding an odd power of energy $E^{L}\cos Lk_{x}(k_{y})$, which is forbidden. Thus, the dominant terms become $\cos 2Lk_{x}$ and $\cos 2Lk_{y}$, whose coefficients are energy-independent, resulting solely from diagonal elements of the determinant, leaving no opportunity for involving $E$-dependent terms. Therefore, transitions associated with extended and localized states do not depend on energy, implying the absence of MEs.
@@
For the critical states, characterized by dispersions of the form $\cos n_{x}Lk_{x}\cos n_{y}Lk_{y}$ with $n_{x},n_{y}\geq 1\in\mathbf{\mathbb{Z}}$, we similarly demonstrate energy independence. The chiral symmetry leads to two distinct scenarios: uniform or purely quasiperiodic hopping matrices $\Pi_{j}$, with $M_{j}$ always purely quasiperiodic (details in Appendix A).
@@
In the first scenario, with unifrom $\Pi_{j}$, the leading dispersion $\cos Lk_{x}\cos Lk_{y}$ emerges from products of exactly $L$ factors of $p_{j}$ (each contributing $e^{ik_{x}}$) and $L$ factors of $m_{j}$ (each contributing $e^{ik_{y}}$), making it energy-independent.
@@
In the second scenario, where both $\Pi_{j}$ and $M_{j}$ are purely quasiperiodic, the leading dispersions include terms such as $\cos 2Lk_{x}\cos Lk_{y}$, $\cos Lk_{x}\cos 2Lk_{y}$, and $\cos 2Lk_{x}\cos 2Lk_{y}$ footnote_dispersion , all of which are energy-independent. The dispersion $\cos Lk_{x}\cos 2Lk_{y}$ arises from product of $L$ factors of $p_{j}$ (each contributing $e^{ik_{x}}e^{ik_{y}}$) and $L$ factors of $m_{j}$ (each contributing $e^{ik_{y}}$), containing exactly $2L$ factors, and thus remains free of energy-dependent contributions. Similarly, the other two dispersions originate from product of $2L$ factors of $p_{j}$ (each contributing $e^{ik_{x}}e^{ik_{y}}$), which generically lead to both $\cos 2Lk_{x}\cos Lk_{y}$ and $\cos 2Lk_{x}\cos 2Lk_{y}$ footnote_rg , both of which are energy-independent. Consequently, critical states also exhibit energy-independent transitions, ensuring no associated MEs appear.
@@
The proof of the criterion highlights the crucial role of chiral symmetry in excluding MEs for the present spin-$1/2$ quasiperiodic systems. It therefore provides general guidelines for realizing pure phases without MEs, as well as coexisting phases with MEs. For coexisting phases, one way to realize is to directly break the condition by adding higher QP frequencies to the system. Alternatively, one may break the chiral symmetry of the system by introducing spin-conserved hopping terms or QP chemical potential into the Hamiltonian footnote_chiral .

### v4

@@ II.3 Renormalization group
The renormalization group (RG) approach (Suslov, 1982; Thouless and Niu, 1983; Ostlund and Pandit, 1984; Niu and Nori, 1986, 1990; Szabó and Schneider, 2018; Jagannathan, 2021; Gonçalves et al., 2022, 2023b) investigates the relevance of renormalized coefficients by iterating the rational approximations of the QP parameter. Specifically, one takes a rational sequence $\alpha^{(n)}=p_{n}/q_{n}$ to define a sequence of periodic Hamiltonians $H^{(n)}$, which approach the quasiperiodic regime only at the irrational limit $\alpha=\alpha^{(\infty)}$. The RG flow to $n\rightarrow\infty$ determines the properties of the states at quasiperiodic regime. With the rational sequence, the energy dispersion of $H^{(n)}$ is a periodic function of quasi-momenta $\kappa_{x}$ and $\kappa_{y}$. To systematically investigate how the relevant coefficients flow as the system size increases, we compute the characteristic determinant $P^{(n)}=|{\cal H}^{(n)}-E|$ at a target energy $E$ for a given size $L=F_{n}$, which is given by
@@
where $\epsilon_{R}^{(n)}$ represents irrelevant higher harmonic terms. The asymptotic nature of eigenstates at $n\rightarrow\infty$ is determined by comparing the relative magnitudes of renormalized hopping coefficients. For extended states, the dominant hopping is along $x$ direction, yielding $\big{|}\mu_{R}^{(n)}/t_{R}^{(n)}\big{|},\big{|}V_{R}^{(n)}/t_{R}^{(n)}\big{|}\rightarrow 0$. For localized states, the dominant hopping is along $y$ direction, leading to $\big{|}t_{R}^{(n)}/V_{R}^{(n)}\big{|},\big{|}\mu_{R}^{(n)}/V_{R}^{(n)}\big{|}\rightarrow 0$. For critical states, all three hopping coefficients are equally relevant, such that $\big{|}\mu_{R}^{(n)}/V_{R}^{(n)}\big{|},\big{|}\mu_{R}^{(n)}/t_{R}^{(n)}\big{|}\geq 1$. With the RG flow we can determine the transitions between extend, localized and critical phases, and the mobility edges as the energy dependent transition points.
@@ II.4 Avila’s global theory
When the Hamiltonian in Eq. (1) can be reduced to an effectively spinless quasiperiodic (QP) chain of dressed particles, the system can be analytically characterized using Avila’s global theory Avila (2015). This theory enables the analytical calculation of the Lyapunov exponent (LE), which is the inverse of the localization length. The method provides a rigorous characterization of extended, localized, and critical states. Specifically, the non-negative LE $\gamma(E)$, quantifies the localization properties of eigenstates with energy $E$. If $\gamma(E)>0$, the state has a finite localization length $\xi(E)=\gamma^{-1}(E)$. If $\gamma(E)=0$, the state is delocalized, with a divergent localization length, corresponding to either an extended or a critical state. Critical states arise when the transfer matrix $T_{j}$ of the delocalized states becomes singular Simon and Spencer (1989); Jitomirskaya and Marx (2012). This singularity can emerge from a universal physical mechanism that the hopping couplings have IDZs Zhou et al. (2023), or the QP potential has incommensurately distributed divergence points, which effectively divides the system into infinite segments, acting as IDZs.
Furthermore, Avila’s global theory can also be employed to derive the analytical expression for the correlation length of extended states. This is achieved by performing the transformation of the system into dual space. Then one can calculate the localization length of localized states in the dual space, which is equivalent to the correlation length of the corresponding extended states in the real space.
Now we outline the procedure for obtaining the LE using Avila’s global theory. For the generic 1D spinful QP Hamiltonian in Eq. (1), we consider the 4-by-4 transfer matrix $T_{j}$ given by $(\begin{array}[]{cccc}u_{j+1\uparrow},&u_{j+1\downarrow},&u_{j\uparrow},&u_{j\downarrow}\end{array})^{\intercal}=T_{j}(\begin{array}[]{cccc}u_{j\uparrow},&u_{j\downarrow},&u_{j-1\uparrow},&u_{j-1\downarrow}\end{array})^{\intercal}$, where $u_{j\sigma}$ is the wave function of eigenstate at site $j$ with spin $\sigma$. In general, the 4-by-4 transfer matrix does not guarantee exact solvability. Exact solvability arises when the transfer matrix is reduced to a block-diagonal 2-by-2 form. As we shall show in Sec. III.3, the present spinful system can be reduced to effective spinless QP model of dressed particles when certain local constraint is introduced. Consequently, the effective transfer matrix becomes a 2-by-2 matrix. In this regime, the LE $\gamma_{\epsilon}(E)$ is computed as
@@
Here, $E$ is the energy of the corresponding eigenstate, $||A||$ is the norm of the matrix A, i.e. the square root of the largest eigenvalue of $A^{\dagger}A$, ${\cal T}_{n,1}=\prod_{j=1}^{n}T_{j}$. The $\epsilon$ is the imaginary part of complexified quasiperiodic phase shift $\theta+i\epsilon$. The LE is generally challenging to compute analytically, even for the simple 2-by-2 transfer matrix $T_{j}$. Nevertheless, a key result from Avila’s global theory is that with the complexification one can analytically compute the LE for all eigenstates, as long as the LE $\gamma_{\epsilon}(E)$ is convex, continuous and piecewise linear with quantized right-derivative, given by
@@
In this case, the nonzero part of derivative of LE is quantized to an integer for all range of $\epsilon$: $d\gamma_{\epsilon}/d\epsilon=2\pi\mathbb{Z}$. This condition also implies that there is no turning point before $\gamma_{\epsilon}$ intersects the $x-$ or $y-$axis, as illustrated in Fig. 1(b). This has an important consequence that the analytic expression for the LE can be determined in the following way: one first computes the complexified LE $\gamma_{\epsilon}$ at the limit $\epsilon\to\infty$, which can be obtained straightforwardly. Then, using the properties that $\gamma_{\epsilon}(E)$ is convex, continuous, and piecewise linear with a quantized right-derivative, the analytic expression of the LE determined at $\epsilon\to\infty$ can be extended to all values of $\epsilon$. Finally, the exact LE is obtained by taking back $\gamma_{\epsilon=0}$, which represents the original physical Lyapunov exponent that characterizes the properties of all eigenstates.
Figure 2: Three universal results for the spinful quasiperiodic chains. (a) Criteria for the system to exhibit pure phases without mobility edges. Here, $\lambda$ is the tuning parameter and $E$ is the energy. States A and B refer to extended, localized, or critical states. The energy-dependent transition points $\lambda_{c}(E)$ become energy-independent, denoted as $\lambda_{c}$, indicates transitions from a regime with mobility edges (MEs) to a pure phase without MEs. (b) Mechanism for the emergence of critical states in the 1D spinful quasiperiodic (QP) system, generated by dual-invariant generalized incommensurate zeros (GIZs) in matrix elements, marked as the blue and red circles in real and dual spaces. (c) Exact solvability of QP systems from local constraint. The local constraint that reduces the generic 1D spinful QP chain into a 1D spinless QP chain of dressed particles (represented by orange spheres), with energy-dependent or energy-independent nearest-neighbor hopping coefficients $t_{j}^{\mathrm{eff}}(E)$ and on-site potentials $V_{j}^{\mathrm{eff}}(E)$, rendering the system exactly solvable through Avila’s global theory.
@@ III Universal results for the quasiperiodic spinful chains
In this section, with the above theoretical approaches, we show three universal results (theorems) for the spinful quasiperiodic systems, which facilitates to establish a unified framework of all the fundamental localization phases, as summarized in Fig. 2. First, we prove a generic condition under which mobility edges (MEs) disappear in the system, namely, the pure phases are obtained under this condition. Second, we uncover a new universal mechanism that critical states are protected by the generalized incommensurate zeros in matrix elements. Third, we outline the conditions, under which the Avila’s global theory can be applied to obtain analytic solutions to the localization properties, correlation length, the associated phase transition points, and mobility edges.
@@ III.1 Criteria for pure phases without MEs
We firsts show that the system hosts only pure phases without MEs under the following conditions: The on-site matrix $M_{j}$ are purely quasiperiodic, and the hopping coupling matrix $\Pi_{j}$ are either uniform or purely quasiperiodic, with the system satisfying the chiral symmetry
@@
We show this criterion by proving that all leading coefficients in Eq.(2) are energy-independent, ensuring that the transitions between different types of states are energy independent, and consequently, the system exhibits pure phases without MEs [Fig.2(a)]. For convenience. we transform $\Pi_{j}^{\prime}$ and $M_{j}^{\prime}$ into the local bases such that the on-site coupling matrix is diagonal, given explicitly as
@@
The proof follows from a careful power counting analysis of the energy dependence in characteristic polynomial $P(E;\kappa_{x},\kappa_{y})$ in Eq. (2). We consider an odd system size $L$ without loss of generality. Due to chiral symmetry, eigenenergies appear in pairs $(-E,E)$, and therefore $P(E)$ is an even function of energy $E$. In the power counting analysis, the energy $E$ explicitly contributes a power of $E$, while $p_{j}^{\perp}$ and $p_{j}^{z}$ contribute dispersions of $e^{ik_{x}}$, and $m_{j}$ contributes dispersions of $e^{ik_{y}}$ (with the mapping to 2D system). Here $\kappa_{x}=Lk_{x}$ and $\kappa_{y}=Lk_{y}$. We utilize the periodic structure to simplify the power counting. Thus, the characteristic polynomial can be expressed as the determinant of a $2L\times 2L$ matrix $|A(E)|$ (details in Appendix B), with elements
@@
and $t_{j}^{\pm}=p_{j}^{z}\pm ip_{j}^{\perp}$. This structure ensures the energy terms and momentum dispersions decouple naturally, eliminating cross terms such as $Ee^{ik_{x}}$. Below, we analyze separately the dominant dispersions for extended, localized, and critical states.
@@
For the extended and localized states, characterized by dispersions of the form $\cos n_{x}Lk_{x}$ and $\cos n_{y}Lk_{y}$, respectively, where $n_{x},n_{y}\geq 1\in\mathbb{Z}$, we show that the dominating coefficients are energy-independent. Specifically, the leading dispersion $\cos Lk_{x}$ and $\cos Lk_{y}$ vanish due to chiral symmetry, as these terms arise from the product of $L$ factors of $e^{ik_{x}}$ and $L$ factors of energy $E$, yielding an odd power of energy $E^{L}\cos Lk_{x}(k_{y})$, which is forbidden. Thus, the dominant terms become $\cos 2Lk_{x}$ and $\cos 2Lk_{y}$, whose coefficients are energy-independent, solely given by diagonal elements of the determinant, leaving no opportunity to involve $E$-dependent terms. Therefore, transitions for extended and localized states do not depend on energy, implying the absence of MEs.
@@
For the critical states, characterized by dispersions of the form $\cos n_{x}Lk_{x}\cos n_{y}Lk_{y}$ with $n_{x},n_{y}\geq 1\in\mathbf{\mathbb{Z}}$, we show below that they are energy independent. We consider two distinct scenarios for the chiral symmetry: uniform and purely quasiperiodic hopping matrices $\Pi_{j}$, with $M_{j}$ always purely quasiperiodic (details in Appendix B).
@@
In the first scenario, for the uniform $\Pi_{j}$, the leading dispersion $\cos Lk_{x}\cos Lk_{y}$ emerges from products of exactly $L$ factors of $p_{j}$ (each contributing $e^{ik_{x}}$) and $L$ factors of $m_{j}$ (each contributing $e^{ik_{y}}$). Thus the momentum terms are exhausted, leaving no momentum dependence of the coefficients, namely, being energy-independent.
@@
In the second scenario, where both $\Pi_{j}$ and $M_{j}$ are purely quasiperiodic, the leading dispersions include terms such as $\cos 2Lk_{x}\cos Lk_{y}$, $\cos Lk_{x}\cos 2Lk_{y}$, and $\cos 2Lk_{x}\cos 2Lk_{y}$ footnote_dispersion , all of which are energy-independent. The dispersion $\cos Lk_{x}\cos 2Lk_{y}$ arises from product of $L$ factors of $p_{j}$ (each contributing $e^{ik_{x}}e^{ik_{y}}$) and $L$ factors of $m_{j}$ (each contributing $e^{ik_{y}}$), containing exactly $2L$ factors, and thus remains free of energy-dependent contributions. Similarly, the other two dispersions originate from product of $2L$ factors of $p_{j}$ (each contributing $e^{ik_{x}}e^{ik_{y}}$), which generically lead to both $\cos 2Lk_{x}\cos Lk_{y}$ and $\cos 2Lk_{x}\cos 2Lk_{y}$ footnote_rg , whose coefficients are both energy-independent. Consequently, critical states also exhibit energy-independent transitions, eliminating the presence of MEs.
@@
The proof of the criterion highlights the crucial role of chiral symmetry in excluding MEs for the present spinful quasiperiodic systems. It therefore provides general guidelines for realizing pure phases without MEs, as well as coexisting phases with MEs. For coexisting phases, one way to realize is to directly break the condition by adding higher QP frequencies to the system. Alternatively, one may break the chiral symmetry of the system by introducing spin-conserved hopping terms or QP chemical potential into the Hamiltonian footnote_chiral .